# Jinhe Bi1,2,4  Danqi Yan1  Yifan Wang1  Aniri1  Wenke Huang3  Zengjie Jin1  Xiaowen Ma1 Artur Hecker2Mang Ye3   Xun Xiao2Hinrich Schütze1,4 Volker Tresp1,4   Yunpu Ma1,411footnotemark: 1

### v2

@@ Appendix D Limitations of Existing Methods and Our Geometric Motivation
D.1 A Critical Review of Existing Selection Paradigms
@@
2. Prohibitive Computational Cost. The second major failure mode is excessive computational overhead, particularly in Training-Based methods. For instance, techniques like TIVE (Liu et al., 2024d) require a costly "warm-up" phase, often involving LoRA-based (Hu et al., 2021) fine-tuning, simply to generate the gradient signals needed for selection. As we show in our runtime analysis (Table 7), the time spent on selection alone can exceed the time saved during the subsequent tuning, violating Principle 2 (Net Efficiency Gain) and rendering the approach impractical at scale.
@@
D.2 The Unifying Cause: A Deep Dive into Representation Anisotropy
We argue that these seemingly disparate issues are symptoms of a single, deeper problem: a failure to account for the pathological geometry of MLLM feature spaces. In practice, high-dimensional embeddings learned by deep networks are rarely isotropic (uniformly distributed); instead, they are highly anisotropic, exhibiting a strong directional bias that corrupts the very notion of semantic distance and plagues models across modalities (Ethayarajh, 2019; Gu et al., 2019; Jing et al., 2021).
@@ D.2 The Unifying Cause: A Deep Dive into Representation Anisotropy
Appendix E Theoretical Foundation: Anisotropy as a Source of Computational Overhead
@@
E.1 The Geometric Pathology of Anisotropy
@@
E.2 The Inevitable Complexity of Anisotropic Selection
@@
Appendix F Benchmark Details
@@
Appendix G Detailed Runtime Analysis

### v3

@@ Appendix E Limitations of Existing Methods and Our Geometric Motivation
E.1 A Critical Review of Existing Selection Paradigms
@@
2. Prohibitive Computational Cost. The second major failure mode is excessive computational overhead, particularly in Training-Based methods. For instance, techniques like TIVE (Liu et al., 2024d) require a costly "warm-up" phase, often involving LoRA-based (Hu et al., 2021; Bi et al., 2025a) fine-tuning, simply to generate the gradient signals needed for selection. As we show in our runtime analysis (Table 13), the time spent on selection alone can exceed the time saved during the subsequent tuning, violating Principle 2 (Net Efficiency Gain) and rendering the approach impractical at scale.
@@
E.2 The Unifying Cause: A Deep Dive into Representation Anisotropy
We argue that these seemingly disparate issues are symptoms of a single, deeper problem: a failure to account for the pathological geometry of MLLM feature spaces. In practice, high-dimensional embeddings learned by deep networks are rarely isotropic (uniformly distributed); instead, they are highly anisotropic, exhibiting a strong directional bias that corrupts the very notion of semantic distance and plagues models across modalities (Ethayarajh, 2019; Gu et al., 2019; Jing et al., 2021; Jiang et al., 2025b, a; Chen et al., 2025b, a; Zhang et al., 2023; Rong et al., 2025; Wang et al., 2025).
@@ E.2 The Unifying Cause: A Deep Dive into Representation Anisotropy
Appendix F Theoretical Foundation: Anisotropy as a Source of Computational Overhead
@@
F.1 The Geometric Pathology of Anisotropy
@@
F.2 The Inevitable Complexity of Anisotropic Selection
@@
Appendix G Cross-Architecture Diagnosis of Representation Anisotropy
To further validate that representation anisotropy is a universal characteristic of Multimodal Large Language Models (MLLMs), we expand our diagnostic analysis across five distinct model-encoder combinations. This section provides detailed statistical distributions of per-dimension means to characterize the geometric shift in the feature space.
G.1 Metric Definitions
We utilize the following statistics to quantify the distribution of per-dimension means across different datasets:
  • Min / Max: The extreme values of the per-dimension mean vectors.
  • Mean($|x|$): The arithmetic mean of the absolute per-dimension means, serving as a primary indicator of the magnitude of the global shift from the origin.
  • Percentiles (P25, P75, P99): Values representing the spread and skewness of the mean distribution across dimensions.
G.2 Per-dimension Mean Statistics across Configurations
Tables 8 through 11 present the statistics for diverse language backbones and vision encoders. Across all configurations, a consistent pattern emerges: textual features remain well-centered around zero ($\text{Mean}(|x|)\approx 0.11\text{--}0.13$), whereas all visual instruction datasets exhibit substantial non-zero shifts ($\text{Mean}(|x|)\approx 0.36\text{--}0.66$).
Table 8: Statistics for Vicuna v1.5 7B + CLIP ViT-L/14 336px.
Dataset Min P25 P75 P99 Max Mean($|x|$)
Text -2.8851 -0.1033 0.0064 0.1105 0.3956 0.1310
A-OKVQA -4.0186 -0.3351 0.0052 0.3329 1.2545 0.4045
GQA -4.0984 -0.3448 0.0057 0.3404 1.2815 0.4141
LLaVA -4.0885 -0.3427 0.0046 0.3394 1.2819 0.4126
OCRVQA -5.6103 -0.5129 0.0023 0.5187 1.8318 0.6127
VG -4.1034 -0.3449 0.0062 0.3413 1.2827 0.4147
Table 9: Statistics for Phi-2 2.7B + SigLIP-SO400M-Patch14-384.
Dataset Min P25 P75 P99 Max Mean($|x|$)
Text -2.9124 -0.1047 0.0072 0.1119 0.4012 0.1328
A-OKVQA -4.3215 -0.3587 0.0048 0.3612 1.3421 0.4384
GQA -4.4102 -0.3701 0.0053 0.3689 1.3745 0.4501
LLaVA -4.4018 -0.3675 0.0042 0.3668 1.3751 0.4483
OCRVQA -6.0321 -0.5514 0.0019 0.5589 1.9647 0.6589
VG -4.4147 -0.3704 0.0058 0.3695 1.3754 0.4505
Table 10: Statistics for Qwen2.5 7B Base + CLIP ViT-L/14 336px.
Dataset Min P25 P75 P99 Max Mean($|x|$)
Text -2.6734 -0.0891 0.0058 0.0945 0.3512 0.1089
A-OKVQA -3.7845 -0.3012 0.0047 0.3089 1.1523 0.3698
GQA -3.8621 -0.3101 0.0052 0.3154 1.1789 0.3789
LLaVA -3.8534 -0.3079 0.0041 0.3147 1.1793 0.3776
OCRVQA -5.1987 -0.4687 0.0021 0.4823 1.6845 0.5634
VG -3.8665 -0.3104 0.0056 0.3162 1.1795 0.3793
Table 11: Statistics for Llama-3-8B + CLIP ViT-L/14 336px.
Dataset Min P25 P75 P99 Max Mean($|x|$)
Text -2.6512 -0.0876 0.0055 0.0929 0.3445 0.1067
A-OKVQA -3.7621 -0.2989 0.0046 0.3067 1.1434 0.3667
GQA -3.8389 -0.3078 0.0051 0.3132 1.1701 0.3756
LLaVA -3.8301 -0.3056 0.0040 0.3125 1.1705 0.3743
OCRVQA -5.1623 -0.4634 0.0020 0.4789 1.6712 0.5589
VG -3.8432 -0.3081 0.0055 0.3140 1.1707 0.3760
G.3 Architectural Generality and Observations
The results demonstrate that representation anisotropy is a robust and persistent phenomenon across different language backbones (Vicuna, Phi-2, Qwen, Llama) and vision encoders (CLIP, SigLIP). While more advanced language models slightly reduce the absolute magnitude of visual anisotropy (e.g., Llama-3 and Qwen show approximately 10% lower $\text{Mean}(|x|)$ than Vicuna), the visual-textual gap remains pronounced in all cases. This evidence confirms that the geometric origins of Global Semantic Drift are general to the MLLM architecture and are not specific to the LLaVA framework.
Appendix H Benchmark Details
@@
Appendix I Sensitivity Analysis of the Selection Ratio
In this section, we provide a detailed sensitivity analysis of the selection ratio $\tau$, which governs the proportion of data retained for instruction tuning. Unlike static hyperparameters, $\tau$ serves as a flexible selection budget that practitioners can adjust based on specific computational constraints. Table 12 presents the performance trade-offs across various sampling ratios. Our analysis reveals that PRISM is robust across diverse selection budgets:
  • Monotonic Visual Scaling: Multimodal performance scales consistently with the data budget, with the 30% ratio (PRISM-Instruct-250K) effectively surpassing the full-dataset baseline.
  • Language Capability Preservation: Textual performance remains remarkably high even at extreme pruning levels (e.g., 5%), suggesting that PRISM’s redundancy identification excels at isolating a high-quality "data core" that minimizes catastrophic forgetting.
Sampling Ratio ($\tau$) Visual Performance Textual Performance
5% 92.3 104.4
10% 95.1 102.9
20% 97.1 102.1
[Uncaptioned image] 30% (PRISM Default) 101.7 101.9
Full Fine-tuning (Baseline) 100.0 100.0
Table 12: Performance robustness across varied selection ratios $\tau$. Visual performance denotes the relative average score across multimodal benchmarks, while textual performance reflects foundational language understanding scores relative to the baseline.
Appendix J Selection Metric Rationale: Correlation vs. Second-Order Distance
The design of PRISM prioritizes a training-free, computationally efficient pipeline. While second-order measures like Mahalanobis distance are theoretically relevant to anisotropy, they introduce significant bottlenecks in large-scale multimodal settings. We justify the use of implicit re-centering via Pearson correlation based on the following factors:
J.1 Algorithmic Efficiency and Scalability
The estimation of Mahalanobis distance requires the computation and inversion of a covariance matrix $\Sigma\in\mathbb{R}^{d\times d}$. For modern MLLMs with high-dimensional feature spaces, this incurs an $\mathcal{O}(Nd^{2})$ cost for estimation and $\mathcal{O}(d^{3})$ for inversion. In contrast, PRISM’s correlation-based scoring operates with near-linear complexity relative to the feature dimension, enabling the processing of the LLaVA-665K dataset in just 1.5 GPU hours.
J.2 Numerical Stability in Anisotropic Spaces
Our singular value analysis demonstrates that pre-trained visual features occupy a low effective rank, where variance is concentrated in a few dominant dimensions. This geometric degeneration often results in ill-conditioned covariance matrices. Relying on matrix inversion in such spaces would require additional regularization (e.g., eigenvalue clipping or shrinkage), introducing new hyperparameters and compromising the simplicity of the training-free paradigm.
Appendix K Detailed Runtime Analysis